# A Mixed-Methods Study of Open-Source Software Maintainers On Vulnerability Management and Platform Security Features

### v1

@@ 5.3 Challenges with platform security features
Not enough automation The most listed PSF challenge described is that there is not enough automation ($L=39,I=11$). As described earlier in Section 5.2, OSS maintainers have limited time and resources to address every reported vulnerability quickly. Automation for security-oriented tasks helps promote easy security without additional overhead.
@@
Too much noise and vulnerability scoring issues Listing participants mentioned that too much noise from PSFs ($L=24,I=10$) and issues from vulnerability scores ($L=21,I=9$), e.g., inaccuracy, are particularly challenging to deal with. Participants describe noise as many false alarms, different from spam or low-quality reports, and how noise can take time away from general OSS tasks and in some cases, overwhelm or scare OSS maintainers ($L=4,I=6$).
@@
Absent CI processes in private forks, failing tests and broken builds in the public eye The next most listed challenge with PSFs is the lack of CI processes when developing fixes on a private fork ($L=8,I=7$), i.e., as a result of adopting the built-in private vulnerability reporting feature. Although private forks are a recommended method of working on fixes with reporters, not being able to run CI processes on such fixes in the same environment as merging pull requests on the main branch can be problematic, resulting in overhead.
@@
Cluttered notifications, setup and usage is too manual The least listed PSF challenges are cluttered notifications ($L=4,I=4$) and too much manual effort ($L=3,I=3$), i.e., usability challenges. Interviewees describe that the best security is “easy security” ($I=6$), including the ease of use for PSFs. When security features are user-friendly and require minimal effort to manage, they can be effectively utilized by OSS maintainers without additional overhead.
@@ 5.4 Barriers hindering adoption of platform security features
Lack of awareness and bad UI presentation The most listed PSF barrier is lack of awareness ($L=33,I=12$). PSFs are available for usage so that OSS projects can have robust vulnerability management processes and proper disclosure upon vulnerability patching. This lack of awareness implies that although tools and processes are available to help promote a thriving secure OSS ecosystem, they are underutilized as indicated in prior work measuring OSS PSFs’ usage [8].
@@
Complex to set up or use, not sure what they do Listing participants mentioned that PSFs are too complex to set up or use ($L=26,I=7$). Though not nearly listed as much, participants are “not sure what they do in the first place” ($L=4,I=2$). These perceptions reflect a barrier of adoption since OSS maintainers question their overhead and purpose. Interviewees express a sense of excessive overhead with setting up PSFs ($I=4$). Others cite general ignorance as to why they are not aware of what particular PSFs do ($L=3,I=7$). As a result, they did not bother looking into PSFs. If users are unclear about what these features do or how they are able to benefit their projects, they are less likely to invest effort in setting them up, especially if they are perceived as complex.
@@
Perception that they are unnecessary The next most listed PSF adoption barrier is the perception that they are unnecessary ($L=23,I=6$). When OSS maintainers perceive PSFs as not adding value to their current processes, they will likely refrain from investing additional time and resources into learning about or implementing these tools; thus, undermining recommended efforts to promote a secure OSS ecosystem.
@@
Project reputation and a lack of motivation The remaining barriers to PSF adoption OSS maintainers mention are concerns with project reputation ($L=11,I=4$) and a lack of motivation ($L=5,I=4$). Concerns about maintaining a positive public perception can deter OSS maintainers from implementing PSFs, that would otherwise enhance security, while those without motivation to adopt PSFs may deprioritize or ignore the use of PSFs. The attitude concerning reputation is not unique to maintainers. Alomar et al. [5] observed this attitude at the organizational level, where organizations try to lower the severity ratings of vulnerabilities in their products.
Participants point out fear of negative project reputation as a reason to not adopt PSFs ($L=3,I=3$), e.g., owning a project with previous high or critical CVE-assigned vulnerabilities ($I=1$). On the other hand, some interviewees consider a project without vulnerabilities as suspicious ($I=2$), i.e., patched vulnerabilities are a good sign of a “healthy project” (P14). One interviewee reflected on such dilemma:
@@ 5.5 Opportunities for improvement & support
Assisted vulnerability analysis and triaging Participants expressed the most need for assisted vulnerability analysis and triaging ($L=39,I=9$), i.e., via automation mechanisms. In particular, participants felt that this would be useful for understanding vulnerabilities and their actual impact on OSS projects ($L=34,I=7$) to reduce false positives and noise described in Section 5.3, e.g., “I hope a tool can take a context into consideration and it should help tooling avoid the false positives, and to trigger only this stuff when [vulnerable]” (P7). Further, participants express a need for assistance with developing fixes and generating tests ($L=21,I=6$), which could streamline the remediation process and ensure more effective vulnerability management, e.g., “generate, you know, tests with test cases or something like that, maybe as a way to lighten the burden of actually developing tests” (P5).
@@
PSF setup assistance, checklists, and nudges Participants also showed interest in having assisted PSF setup ($L=37,I=10$). In particular, interviewees mention needs for assistance generating security policy content using project scope ($I=3$), security tooling recommendations based on project contents ($I=6$), and interactive guidance throughout the setup process ($I=7$), e.g., “it’s a little bit not easy to use […] now I need to figure out, okay, what is the proper setup for my project [since] I’m of the mindset that brevity is king.” (P9). Further, participants mention wanting checklists and suggestions of things to do for recommended PSFs ($L=15,I=3$), alongside nudges or reminders and “easy” configurations as methods of encouraging others to strengthen their projects’ security posture ($L=16,I=5$). One interview elaborates on their view of overall PSFs: “It’s pretty overwhelming, and you might just be like, why do I need any of these? Like, why is this important? Having a single button that just says, enable best practices, would go a long way.” (P16).
@@
Security funding and cyber defense gamification The next most-listed want for participants is security-specific funding ($L=32,I=7$), i.e., funding to be used for security-related tasks and efforts. In particular, participants mention using funds for maintainers and reporters to get paid ($L=18,I=6$), funding a dedicated bounty pool for valid vulnerabilities ($I=3$), or requesting annual reviews from security experts ($I=2$). There was also interest in cyber defense gamification ($L=13,I=6$), i.e., to recognize projects that adopt vulnerability management processes. Four interviewees mention how they use reporter reputation to determine vulnerability report legitimacy. Such participants and others expressed interest in project reputation icons, e.g., a “green shield” ($I=3$), for projects that adopt recommended PSFs and tooling ($I=7$); thus, recognizing and rewarding projects that exhibit a proper security posture. “Everyone wants to have the green shield […] This community health status people see, stuff like that, like you’ll be surprised by how many people want to feel those bars.” (P10).
@@
User-friendly resources and documentation Lastly, participants expressed interest in having more user-friendly documentation and resources ($L=17,I=8$), some citing general “ignorance” of proper OSS security ($I=6$). In particular, participants mention OSS recommended training material for “normie developers” ($L=3,I=4$), e.g., free webinars ($I=2$), redirection to related projects with implemented security practices ($L=7,I=3$), and “easier directions” for setting up and what to expect from PSFs ($I=6$), e.g., not being “scared” of false positives ($I=3$). “Security best practices and tooling and all of that is tribal knowledge, and it shouldn’t be.” (P12).
@@ 6 Discussion
In this section, we present implications for OSS platforms and OSS maintainers to alleviate prominent challenges we discovered, explore areas for reducing PSF barriers, and expand the benefits participants expressed as helpful for vulnerability management. Further, we present potential research directions aimed at supporting OSS project maintainers during the vulnerability management lifecycle.
@@ 6.1 Implications & future directions based on collective challenges
When maintainers wait for upstream dependencies to fix a vulnerability, there is a delay in addressing security issues within their projects, thus exposing OSS projects to potential exploits. Tools that automate identifying and prioritizing vulnerabilities can help maintainers quickly determine which issues require immediate attention and which can be deferred. Tooling should be explored that implements mechanisms for automatically identifying the impact of vulnerabilities using source code as context, i.e., to reduce noise from security tooling. To that end, future research can leverage LLMs to (1) help OSS maintainers with interpreting reported vulnerabilities, e.g., by leveraging OSS security datasets like those curated by OpenSSF [38], and (2) help OSS maintainers generate patches for reported vulnerabilities with minimized regression tests by using the contents of disclosed reports. OSS platforms should consider providing CI feature capabilities in private forks so OSS maintainers can expedite fixing vulnerabilities and reduce costs of regression tests, which can be very high [23, 53], indicating a need for research focused on regression testing for security. GitHub disallows private forks from using CI features for security purposes [41], but OSS maintainers desire such features in our studies, especially for fixing vulnerabilities, suggesting the need for further research on secure CI features for vulnerability management.
@@ 6.2 Implications & future directions based on PSF adoption barriers
We encourage OSS maintainers to enable PSFs, e.g., especially those that are togglable by nature such as private vulnerability reporting, and recommend OSS platforms provide an Enable best practices button for PSFs that require no-to-minimal setup and include links to respective PSFs that are user-friendly for OSS maintainers without security backgrounds. We discourage OSS maintainers from being complacent with users submitting vulnerabilities using public GitHub issues—which is already discouraged [40]—and should instead take advantage of the Private vulnerability reporting PSF [85] regardless of project size, a PSF we recommend OSS platforms enable by default.
OSS platforms and researchers should also consider working towards functionalities that (1) improve the quality and presentation of security notifications; (2) automatically inform contributors about if their issue should be filed as a vulnerability report based on past OSS vulnerabilities; and (3) features to reward projects with recommended security posture and be designed in such a way that creates a positive reputation. Research in such directions can focus on creating better-designed PSFs to provide more value to OSS maintainers, such that PSFs are easier to use, provide guidance to reporters, and promote positivity with their adoption.
@@ 8 Ethical considerations
Appendix A Listing survey questions and outline
[Study information, eligibility criteria, privacy notice, etc.]
Eligibility question Do you reside in an OFAC-sanctioned region or are affiliated with an OFAC-sanction entity?
Main questionnaire:
1. How do you currently handle vulnerability reports in your OSS project(s)? Please describe in detail.
2. What tooling or suggested practices, if any, do you use to track and manage vulnerabilities? Please list any software, services, or methodologies you rely on.
3. How often do you review and update, if necessary, your vulnerability management practices? Please describe in detail.
4. What are the challenges you currently face in managing OSS vulnerabilities? Please list all that you can think of.
5. How would you describe your overall experience(s) with OSS vulnerability management? Please describe in detail.
6. Which platform security features do you currently have enabled? Please list all that you have enabled.
7. What challenges have you faced maintaining or publishing a security policy? Please list all that you can think of.
8. Do you have any suggestions for policies or guidelines that OSS platforms should implement to improve vulnerability management across repositories? If so, please list any policy or guideline recommendations.
9. What challenges have you faced maintaining and publishing security advisories? Please list all that you can think of.
10. What challenges have you faced when considering or using the Private vulnerability reporting feature? Please list all that you can think of.
11. If applicable, please list any alternative mechanisms of reporting (e.g., email) or disclosure you use outside of platform security features. Please list all that you can think of.
12. What resources or tools you have found useful for managing vulnerabilities? Please list all that you can think of.
13. Do you feel that the platform security tools provided by OSS platforms are sufficient for managing vulnerabilities? If not, what additional features or improvements would you like to see? Please describe in detail.
14. What resources or support would be most beneficial for you to improve your vulnerability management practices? Please list all that you can think of.
[Background multiple choice questions (Table 1)]
Appendix B Semi-structured interview guide
0. Project maintainer duties
0. Can you tell me about your duties as an open-source project maintainer?
1. Current vulnerability management practices
1. Can you elaborate on your process for handling OSS vulnerabilities?
2. What tools or platforms do you find the most effective for tracking and managing vulnerabilities?
3. Are there any aspects of these tools that you find particularly helpful?
4. What tools or platforms do you find the least effective for tracking and managing vulnerabilities?
5. Are there any features of these tools that you find particularly lacking?
6. How do you decide when to review and update your vulnerability management practices? Can you provide an example of a recent update and what prompted it?
2. Challenges with vulnerability management
7. In the survey, you mentioned several challenges in managing vulnerabilities, including [mention responses]. Can you provide more details about the most significant challenge(s)?
8. How have these challenges impacted your project and overall workflow?
9. Are there any particular hurdles you overcame? How?
10. How do these hurdles affect your ability to manage vulnerabilities effectively? Can you provide an example?
3. Challenges with platform security features
11. In the survey, you mentioned several challenges when using specific platform security features, including [mention responses]. Can you provide more details about the most significant challenge(s)?
12. How have these challenges impacted your project and overall workflow?
13. Are there any particular hurdles you overcame? How?
14. How do these hurdles affect your ability to manage vulnerabilities effectively? Can you provide an example?
4. Barriers to adopting platform security features
15. In the survey, you mentioned several barriers when considering using specific platform security features, including [mention responses]. Can you provide more details about the most significant barrier(s)?
16. I noticed you did not mention using [platform security feature(s)]. Could you elaborate?
17. What do you think would motivate you, if possible, to adopt more platform security features?
5. Opportunities for improvement and support
18. You mentioned needing [resources and support] for vulnerability management. Can you expand on the specific types of support that would be most beneficial?
19. In your opinion, what additional features or improvements should OSS platforms implement to better support vulnerability management?
20. How would these improvements make a difference in your project? Workflow?
21. How would you see emerging technologies, such as large language models, support you?
22. Is there anything else you would like to add about software vulnerability management in your OSS project(s)?
Appendix C Interview participants
Appendix D Full table of curated codes

### v2

@@ 5.3 Challenges with platform security features
Not enough automation (F17) The most listed PSF challenge described is that there is not enough automation ($L=39,I=11$). As described earlier in Section 5.2, OSS maintainers have limited time and resources to address every reported vulnerability quickly. Automation for security-oriented tasks helps promote easy security without additional overhead.
@@
Too much noise and vulnerability scoring issues (F18 & F19) Listing participants mentioned that too much noise from PSFs ($L=24,I=10$) and issues from vulnerability scores ($L=21,I=9$), e.g., inaccuracy, are particularly challenging to deal with. Participants describe noise as many false alarms, different from spam or low-quality reports, and how noise can take time away from general OSS tasks and in some cases, overwhelm or scare OSS maintainers ($L=4,I=6$).
@@
Absent CI processes in private forks, failing tests and broken builds in the public eye (F20 & F21) The next most listed challenge with PSFs is the lack of CI processes when developing fixes on a private fork ($L=8,I=7$), i.e., as a result of adopting the built-in private vulnerability reporting feature. Although private forks are recommended for working on fixes with reporters, not being able to run CI processes on such fixes in the same environment as merging pull requests on the main branch can be problematic, resulting in overhead.
@@
Cluttered notifications, setup and usage is too manual (F22 & F23) The least listed PSF challenges are cluttered notifications ($L=4,I=4$) and too much manual effort ($L=3,I=3$), i.e., usability challenges. Interviewees describe that the best security is “easy security” ($I=6$), including the ease of use for PSFs. When security features are user-friendly and require minimal effort to manage, they can be effectively utilized by OSS maintainers without additional overhead.
@@ 5.4 Barriers hindering adoption of platform security features
Lack of awareness and bad UI presentation (F24 & F28) The most listed PSF barrier is lack of awareness ($L=33,I=12$). PSFs are available for usage so that OSS projects can have robust vulnerability management processes and proper disclosure upon vulnerability patching. This lack of awareness implies that although tools and processes are available to help promote a thriving secure OSS ecosystem, they are underutilized as indicated in prior work [11].
@@
Complex to set up or use, not sure what they do (F25 & F30) Listing participants mentioned that PSFs are too complex to set up or use ($L=26,I=7$). Though not nearly listed as much, participants are “not sure what they do in the first place” ($L=4,I=2$). These perceptions reflect a barrier of adoption since OSS maintainers question their overhead and purpose. Interviewees express a sense of excessive overhead with setting up PSFs ($I=4$). Others cite general ignorance as to why they are not aware of what particular PSFs do ($L=3,I=7$). As a result, they did not bother using PSFs.If users are unclear about what these features do or how they can benefit their projects, they are less likely to invest effort in setting them up, especially if they are perceived as complex.
@@
Perception that they are unnecessary (F26) The next most listed PSF adoption barrier is the perception that they are unnecessary ($L=23,I=6$). When OSS maintainers perceive PSFs as not adding value to their current processes, they will likely refrain from investing additional time and resources into learning about or implementing these tools; thus, undermining recommended efforts to promote a secure OSS ecosystem.
@@
Project reputation and a lack of motivation (F27 & F29) The remaining barriers to PSF adoption OSS maintainers mention are concerns with project reputation ($L=11,I=4$) and a lack of motivation ($L=5,I=4$). Concerns about maintaining a positive public perception can deter OSS maintainers from implementing PSFs, that would otherwise enhance security, while those without motivation to adopt PSFs may deprioritize or ignore the use of PSFs. The attitude concerning reputation is not unique to maintainers. Alomar et al. [9] observed this attitude at the organizational level, where organizations try to lower the severity ratings of vulnerabilities in their products.
Participants point out fear of negative project reputation as a reason to not adopt PSFs ($L=3,I=3$), e.g., owning a project with previous high or critical CVE-assigned vulnerabilities ($I=1$). On the other hand, some interviewees consider a project without vulnerabilities as suspicious ($I=2$), i.e., patched vulnerabilities are a good sign of a “healthy project” (P14). One interviewee reflected on such a dilemma:
@@ 5.5 Opportunities for improvement & support
Assisted vulnerability analysis and triaging (F31) Participants expressed the most need for assisted vulnerability analysis and triaging ($L=39,I=9$), i.e., via automation mechanisms. In particular, participants felt that this would be useful for understanding vulnerabilities and their actual impact on OSS projects ($L=34,I=7$) to reduce false positives and noise described in Section 5.3, e.g., “I hope a tool can take a context into consideration and it should help tooling avoid the false positives, and to trigger only this stuff when [vulnerable]” (P7). Further, participants express a need for assistance with developing fixes and generating tests ($L=21,I=6$), which could streamline the remediation process and ensure more effective vulnerability management, e.g., “generate, you know, tests with test cases or something like that, maybe as a way to lighten the burden of actually developing tests” (P5).
@@
PSF setup assistance, checklists, and nudges (F32, F36, & F37) Participants also showed interest in having assisted PSF setup ($L=37,I=10$). In particular, interviewees mention needs for assistance generating security policy content using project scope ($I=3$), security tooling recommendations based on project contents ($I=6$), and interactive guidance throughout the setup process ($I=7$), e.g., “it’s a little bit not easy to use […] I need to figure out, okay, what is the proper setup for my project [since] I’m of the mindset that brevity is king.” (P9). Further, participants mention wanting checklists and suggestions of things to do for recommended PSFs ($L=15,I=3$), alongside nudges or reminders and “easy” configurations as methods of encouraging others to strengthen their projects’ security posture ($L=16,I=5$). One interview elaborates on their view of overall PSFs: “It’s pretty overwhelming, and you might just be like, why do I need any of these? Like, why is this important? Having a single button that just says, enable best practices, would go a long way.” (P16).
@@
Security funding and cyber defense gamification (F33 & F35) The next most-listed want for participants is security-specific funding ($L=32,I=7$), i.e., funding to be used for security-related tasks and efforts. In particular, participants mention using funds for maintainers and reporters to get paid ($L=18,I=6$), funding a dedicated bounty pool for valid vulnerabilities ($I=3$), or requesting annual reviews from security experts ($I=2$). There was also interest in cyber defense gamification ($L=13,I=6$), i.e., to recognize projects that adopt vulnerability management processes. Four interviewees mention how they use reporter reputation to determine vulnerability report legitimacy. Such participants and others expressed interest in project reputation icons, e.g., a “green shield” ($I=3$), for projects that adopt recommended PSFs and tooling ($I=7$); thus, recognizing and rewarding projects that exhibit a proper security posture. “Everyone wants to have the green shield […] This community health status people see, stuff like that, like you’ll be surprised by how many people want to feel those bars.” (P10).
@@
User-friendly resources and documentation (F34) Lastly, participants expressed interest in having more user-friendly documentation and resources ($L=17,I=8$), some citing general “ignorance” of proper OSS security ($I=6$). In particular, participants mention OSS recommended training material for “normie developers” ($L=3,I=4$), e.g., free webinars ($I=2$), redirection to related projects with implemented security practices ($L=7,I=3$), and “easier directions” for setting up and what to expect from PSFs ($I=6$), e.g., not being “scared” of false positives ($I=3$). “Security best practices and tooling and all of that is tribal knowledge, and it shouldn’t be.” (P12).
[…]
@@ 6 Discussion
In this section, we present additional analyses of the challenges identified and presented in Section 5,implications for OSS platforms and OSS maintainers to alleviate prominent challenges we discovered, explore areas for reducing PSF barriers, and expand the benefits participants expressed as helpful for vulnerability management. Further, we present potential research directions aimed at supporting OSS project maintainers during the vulnerability management lifecycle.
6.1 Bridging factors to reveal underlying issues
OSS maintainers described challenges from vulnerability management and their experiences with PSFs. Uncovering underlying issues can help identify areas of need, why particular PSFs may not be adopted, and inform future solutions.
OSS maintainers described challenges that hinder their ability to take necessary actions from PSF output rooted in overwhelming security duties and missing PSF support that would help vulnerability triaging. Dependency trust issues (F9), e.g., too many false positives from upstream dependencies (F18) and cluttered update notifications (F22 & F28), cause fatigue (F29) and a disinclination to continue PSF adoption. Further, maintainers expressed struggle with developing patches (F10, F14, & F31) since core CI/CD processes are missing from some PSFs (F20) that cause broken tests/builds (F21); thus, leading maintainers to deem PSFs as unnecessary (F26), deploy custom behind-the-scenes processes that can cause burnout (F29), and become overwhelmed in the CVE process (F13, F15, & F16). Without any clear prioritization or impact analysis, some PSFs may seem more like distractions, and the lack of trust demonstrated by maintainers is a reinforcing factor in the perception that PSFs are unnecessary, especially if they seem to add little value to the project. Addressing this concern means incorporating features that would be perceived to provide real value, e.g., timely patching, and the use of gamification or funding (F33 & F35) and checklists (F36) can also be used to reframe these tasks as rewarding and meaningful, promoting a positive narrative for security.
OSS maintainers also described challenges that hinder their ability to adopt PSFs rooted in misunderstanding and misinterpretation of PSFs’ configurations, use cases, and purposes. Knowledge gaps (F10), e.g., around vulnerability scoring procedures (F19), hinder effective prioritization and remediation efforts. Resource constraints (F12), including limited time (F11) and lack of automation (F17), exacerbate these difficulties, often leading to delays in addressing vulnerabilities or adopting PSFs; further, the complexity (F25) and lack of awareness (F24) surrounding PSFs create additional barriers. If OSS maintainers do not fully understand the benefits or use cases for PSFs (F30), or find them confusing, they are unlikely to explore or enable them (F37). Training resources that simplify security concepts and integrate clear documentation (F32 & F34) for PSFs could bridge this gap, empowering maintainers to overcome these hurdles and integrate PSFs effectively into their workflow.
@@ 6.2 Implications & future directions based on collective challenges
When maintainers wait for upstream dependencies to fix a vulnerability, there is a delay in addressing security issues within their projects, thus exposing OSS projects to potential exploits. Tools that automate identifying and prioritizing vulnerabilities can help maintainers quickly determine which issues require immediate attention and which can be deferred. To support assisted vulnerability analysis and triaging, tooling should be explored that implements mechanisms for automatically identifying the impact of vulnerabilities using source code as context, i.e., to reduce noise from security tooling. To that end, future research can leverage LLMs to (1) help OSS maintainers with interpreting reported vulnerabilities, e.g., by leveraging OSS security datasets like those curated by OpenSSF [44], and (2) help OSS maintainers generate patches for reported vulnerabilities with minimized regression tests by using the contents of disclosed reports. Although LLMs are useful for tasks like code summarization and generation  [106, 3], uses of LLMs pose challenges. For instance, LLMs may misinterpret security reports or generate incomplete/inaccurate patches, leading to regressions; further, there may be hesitance to trust AI-generated suggestions because LLMs’ decision-making processes are uninterpretable [58].OSS platforms should consider providing CI feature capabilities in private forks so OSS maintainers can expedite fixing vulnerabilities and reduce costs of regression tests, which can be very high [28, 62], indicating a need for research focused on regression testing for security. GitHub disallows private forks from using CI features for security purposes [48], but OSS maintainers desire such features in our studies, especially for fixing vulnerabilities, suggesting the need for further research on secure CI features for vulnerability management. This might involve sandboxing or access controls, e.g., where private forks can use CI features under controlled conditions.
@@ 6.3 Implications & future directions based on PSF adoption barriers
We encourage OSS maintainers to enable PSFs, e.g., especially those that are togglable by nature, such as private vulnerability reporting, and recommend OSS platforms provide an Enable best practices button for PSFs that require no-to-minimal setup and include links to respective PSFs that are user-friendly for OSS maintainers without security backgrounds. Research challenges for enabling such a button include automatically analyzing the context of the OSS project, including its application-specific code and its dependencies. Another research challenge of such a button would be conducting a developer study to determine the best manner in which to incorporate developer preferences into the automation of best PSF practices provided by the button.
Our study further demonstrates a wide variety of OSS maintainers’ views as to the extent to which vulnerabilities should be publicized, especially before a fix is provided. That spectrum includes a preference for using email due to the more inherent privacy compared to public GitHub issues or having to deal with the usability issues involving PSFs. Although GitHub discourages public vulnerability reporting [47] and, hence, encourages using the private vulnerability reporting PSF [96], our study provides little evidence to back up such a recommendation from OSS maintainers’ perspectives.
Nevertheless, preventing premature disclosure of a potentially severe and easily exploitable vulnerability, which may even be particularly difficult to fix, is a potentially strong argument backing GitHub’s recommendation. To that end, feature research can help automatically identify opportunities for beneficial PSF usage. For example, when a user tries to submit a public issue mentioning keywords like “vulnerability” or “security”, OSS platforms could display a warning and automatically nudge maintainers to encourage enabling lightweight PSFs that require minimal effort to implement.
Other future research opportunities encouraging the use of private vulnerability reporting include producing a convincing means or incentive of recommending using PSF features (e.g., by gamifying the PSFs with badges or achievements) or even enforcing the use of PSFs that OSS platform providers recommend. As an example of such enforcement, future research can automatically convert issues submitted to the public that are likely vulnerabilities or may simply be bugs that, as identified through static reachability analysis, are reachable on the project’s attack surface. Such research would go beyond the coarse-grained identification of vulnerabilities in a GitHub project’s dependencies by helping maintainers triage potential vulnerabilities as early as possible with higher confidence (e.g., as provided through a static or dynamic analysis during report submission). Such a future approach would help alleviate the steep learning curve of understanding software security while promoting the adoption of PSF features that can reduce exploitation of publicly-reported OSS vulnerabilities.
OSS platforms and researchers should also consider working towards functionalities that (1) improve the quality and presentation of security notifications; (2) automatically inform contributors about if their issue should be filed as a vulnerability report based on past OSS vulnerabilities; (3) provide PSF setup assistance, e.g., generating a context-aware security policy;and (4) gamification features to reward projects with recommended security posture and be designed in such a way that creates a positive reputation. Research in such directions can focus on creating better-designed PSFs to provide more value to OSS maintainers, such that PSFs are easier to use, provide guidance to reporters, and promote positivity with their adoption. Future research can also investigate how to effectively use gamification to disseminate user-friendly resources on proper OSS security, allowing any maintainers to effortlessly learn security best practices. For example, a project reputation icon, e.g., “green shield”, for enabling best-practices PSFs can come with documentation on why each PSF should be enabled. Further, PSFs can have a leaderboard of many kinds: maintainers who followed security best practices most, reporters who provide the best-rated (e.g., by maintainers) vulnerability-oriented issues, reporters who provide the easiest-to-understand and most convincing fixes (e.g., avoiding regressions while ensuring security), and even OSS projects with the best security-focused documentation. This can further motivate maintainers to engage with security-related gamification features and available security resources.
@@ 8 Ethics considerations
We did not monetarily compensate survey participants because of (1) our funding limits as academic researchers, and (2) similar to other qualitative studies, compensating international participants is logistically difficult. We focused on creating a study that provided intrinsic value to participants through meaningful engagement and insights. Further, (3) we reduced the time cost of completing our survey by allowing participants to skip any free-response questions and encouraged them not to spend more than 15 minutes filling out our survey, and (4) survey participation is voluntary, OSS participants often support academic research out of intrinsic motivation [41, 46, 5] to help improve the security posture of the OSS ecosystem through our study’s results. This approach ensured that participation gives a sense of purpose and contributing to the greater good, i.e., for further securing the OSS ecosystem, despite the lack of monetary compensation for survey participants. However, each of our 22 interviewees was monetarily compensated \$20.
@@
9 Open science
We make our listing study survey questions, semi-structured interview guide, detailed interview participants’ information, and an extended table of codes available in our artifact [2].